# A Bernstein polynomial approach for the estimation of cumulative distribution functions in the presence of missing data

### v1

@@ 7.2 Proofs of the asymptotic properties of $\smash{\widehat{F}_{n,m}}$
Let $y\in[0,1]$ be given. First, note that a Taylor expansion of $1/\hat{\pi}_{i}(\boldsymbol{X}_{1:n})$ around $1/\pi(\boldsymbol{X}_{i})$ yields
$\frac{1}{\hat{\pi}_{i}(\boldsymbol{X}_{1:n})}=\frac{1}{\pi_{i}(\boldsymbol{X}_{i})}-\frac{1}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))+\sum_{j=2}^{\infty}\frac{(-1)^{j}}{\pi_{i}(\boldsymbol{X}_{i})^{j+1}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))^{j}.$
@@
$\widehat{F}_{n,m}(y)$ $=\widetilde{F}_{n,m}(y)-n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$ (7.9)
$\hskip 56.9055pt+n^{-1}\sum_{j=2}^{\infty}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(-1)^{j}\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})^{j+1}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))^{j}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y).$
@@
$\hat{p}(\boldsymbol{X}_{i})=n^{-1}\sum_{k=1}^{n}\mathds{1}_{\{\boldsymbol{X}_{k}=\boldsymbol{X}_{i}\}}$
@@
$\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i})=\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{\hat{p}(\boldsymbol{X}_{i})}.$
@@ A Taylor expansion of $1/\hat{p}(\boldsymbol{X}_{i})$ around $1/p(\boldsymbol{X}_{i})$ yields
$\frac{1}{\hat{p}(\boldsymbol{X}_{i})}$ $=\frac{1}{p(\boldsymbol{X}_{i})}-\frac{1}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))+\sum_{j=2}^{\infty}\frac{(-1)^{j}}{p(\boldsymbol{X}_{i})^{j+1}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{j}.$
We put this expansion in the previous equation and evaluate term by term:
$\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i})$ $=\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}$ (7.10)
$\quad-\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))$
$\quad+n^{-1}\sum_{\ell=2}^{\infty}\sum_{j=1}^{n}\frac{(-1)^{\ell}}{p(\boldsymbol{X}_{i})^{\ell+1}}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{\ell}.$
@@
$\mathsf{E}\left[\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mid\boldsymbol{X}_{1:n}\right]$
$\quad=\mathsf{E}\left[\frac{n^{-1}\sum_{j=1}^{n}\mathsf{E}[(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}]\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\right]=0,$
@@
$\mathsf{Var}\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mid\boldsymbol{X}_{1:n}\right)=\frac{n^{-2}\sum_{j=1}^{n}\mathsf{Var}(\delta_{j}\mid\boldsymbol{X}_{1:n})\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}\ll\frac{n^{-1}}{p_{\min}^{2}},$
where $p_{\min}\equiv\min_{\boldsymbol{x}}p(\boldsymbol{x})>0$ by Assumption (A2). Hence, using the conditional variance formula,
$\mathsf{Var}\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\right)$
$\quad=\mathsf{E}\left[\mathsf{Var}\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mid\boldsymbol{X}_{1:n}\right)\right]+0\ll\frac{n^{-1}}{p_{\min}^{2}},$
@@
$\mathsf{E}\left[\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\right)^{2}\right]\ll\frac{n^{-1}}{p_{\min}^{2}}.$
@@
$\mathsf{E}\left[\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))\right)^{2}\right]$
$\ll\frac{1}{p_{\min}^{4}}\sqrt{\mathsf{E}\left[\left(n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}\right)^{4}\right]}\sqrt{\mathsf{E}\left[(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{4}\right]}$
@@
$\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))=\mathcal{O}_{L^{2}}(n^{-1}),$ (7.11)
with the rate being dependent on $p_{\min}$. Similarly, the residual tail sum satisfies
$n^{-1}\sum_{\ell=2}^{\infty}\sum_{j=1}^{n}\frac{(-1)^{\ell}}{p(\boldsymbol{X}_{i})^{\ell+1}}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{\ell}=\mathcal{O}_{L^{2}}(n^{-1}),$ (7.12)
so that, from (7.10), we have
$\mathsf{E}\left[(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))^{2}\right]=\mathcal{O}(n^{-1}).$ (7.13)
@@
$\mathsf{E}\left[n^{-1}\sum_{j=2}^{\infty}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(-1)^{j}\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})^{j+1}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))^{j}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]=\mathcal{O}(n^{-1}),$ (7.14)
with the rate being dependent on $p_{\min}$ and $\pi_{\min}$.
@@
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$ (7.15)
$\quad=n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$
$\qquad+n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi_{i}(\boldsymbol{X}_{i})}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y).$
@@
$\mathsf{E}\left[(\delta_{i}-\pi(\boldsymbol{X}_{i}))^{2}\mid\boldsymbol{X}_{1:n}\right]=\mathsf{Var}(\delta_{i}\mid\boldsymbol{X}_{1:n})=\pi_{i}(\boldsymbol{X}_{i})(1-\pi_{i}(\boldsymbol{X}_{i})),$
@@
$\quad=\mathsf{E}\left[\delta_{i}-\pi(\boldsymbol{X}_{i})\mid\boldsymbol{X}_{1:n}\right]\mathsf{E}\left[(\delta_{j}-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}\right]\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}=0.$
Hence, recalling that $\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i})=n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi_{i}(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}/\hat{p}(\boldsymbol{X}_{i})$, we deduce
$\mathsf{E}\left[\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}\right]$
$\quad=\mathsf{E}\left[\frac{\mathsf{E}\left[(\delta_{i}-\pi(\boldsymbol{X}_{i}))^{2}\mid\boldsymbol{X}_{1:n}\right]}{n\,\pi_{i}(\boldsymbol{X}_{i})^{2}\hat{p}(\boldsymbol{X}_{i})}\right]+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\mathsf{E}\left[\frac{\mathsf{E}\big[(\delta_{i}-\pi(\boldsymbol{X}_{i}))(\delta_{j}-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}\mid\boldsymbol{X}_{1:n}\big]}{n\,\pi_{i}(\boldsymbol{X}_{i})^{2}\hat{p}(\boldsymbol{X}_{i})}\right]$
$\quad=\mathsf{E}\left[\frac{\pi_{i}(\boldsymbol{X}_{i})(1-\pi_{i}(\boldsymbol{X}_{i}))}{n\,\pi_{i}(\boldsymbol{X}_{i})^{2}\hat{p}(\boldsymbol{X}_{i})}\right]+0$
@@
$\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]$
$\quad=\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\mathsf{E}\left[\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi_{i}(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}\right]\mathsf{E}[\mathds{1}_{\{Y_{i}\leq k/m\}}\mid\boldsymbol{X}_{1:n}]\,b_{m,k}(y)\right]$

### v2

@@ 7.2 Proofs of the asymptotic properties of $\smash{\widehat{F}_{n,m}}$
Let $y\in(0,1)$ be given. First, note that a Taylor expansion of $1/\hat{\pi}_{i}(\boldsymbol{X}_{1:n})$ around $1/\pi(\boldsymbol{X}_{i})$ yields
$\frac{1}{\hat{\pi}_{i}(\boldsymbol{X}_{1:n})}=\frac{1}{\pi(\boldsymbol{X}_{i})}-\frac{1}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))+\sum_{j=2}^{\infty}\frac{(-1)^{j}}{\pi(\boldsymbol{X}_{i})^{j+1}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))^{j}.$
@@
$\widehat{F}_{n,m}(y)$ $=\widetilde{F}_{n,m}(y)-n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$ (7.9)
$\hskip 56.9055pt+n^{-1}\sum_{j=2}^{\infty}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(-1)^{j}\delta_{i}}{\pi(\boldsymbol{X}_{i})^{j+1}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))^{j}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y).$
@@
$\hat{p}(\boldsymbol{x})=n^{-1}\sum_{k=1}^{n}\mathds{1}_{\{\boldsymbol{X}_{k}=\boldsymbol{x}\}}$
@@
$\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i})=\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{\hat{p}(\boldsymbol{X}_{i})}.$
@@ A Taylor expansion of $1/\hat{p}(\boldsymbol{X}_{i})$ around $1/p(\boldsymbol{X}_{i})$ yields
$\frac{1}{\hat{p}(\boldsymbol{X}_{i})}$ $=\frac{1}{p(\boldsymbol{X}_{i})}-\frac{1}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))+\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell}}{p(\boldsymbol{X}_{i})^{\ell+1}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{\ell}.$
We substitute this expansion into the previous equation and evaluate term by term:
$\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i})$ $=\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}$ (7.10)
$\quad-\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))$
$\quad+n^{-1}\sum_{\ell=2}^{\infty}\sum_{j=1}^{n}\frac{(-1)^{\ell}}{p(\boldsymbol{X}_{i})^{\ell+1}}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{\ell}.$
@@
$\mathsf{E}\left[\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mid\boldsymbol{X}_{1:n}\right]$
$\quad=\mathsf{E}\left[\frac{n^{-1}\sum_{j=1}^{n}\mathsf{E}[(\delta_{j}-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}]\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\right]=0,$
@@
$\mathsf{Var}\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mid\boldsymbol{X}_{1:n}\right)=\frac{n^{-2}\sum_{j=1}^{n}\mathsf{Var}(\delta_{j}\mid\boldsymbol{X}_{1:n})\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}\ll\frac{n^{-1}}{p_{\min}^{2}},$
where $p_{\min}=\min_{\boldsymbol{x}}p(\boldsymbol{x})>0$ by Assumption (A2). Hence, using the conditional variance formula,
$\mathsf{Var}\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\right)$
$\quad=\mathsf{E}\left[\mathsf{Var}\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\mid\boldsymbol{X}_{1:n}\right)\right]+0\ll\frac{n^{-1}}{p_{\min}^{2}},$
@@
$\mathsf{E}\left[\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})}\right)^{2}\right]\ll\frac{n^{-1}}{p_{\min}^{2}}.$
@@
$\mathsf{E}\left[\left(\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))\right)^{2}\right]$
$\ll\frac{1}{p_{\min}^{4}}\sqrt{\mathsf{E}\left[\left(n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}\right)^{4}\right]}\sqrt{\mathsf{E}\left[(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{4}\right]}$
@@
$\frac{n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}}{p(\boldsymbol{X}_{i})^{2}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))=\mathcal{O}_{L^{2}}(n^{-1}),$ (7.11)
where the constant implicit in the error term depends on $p_{\min}$. Similarly, the residual tail sum in (7.10) satisfies
$n^{-1}\sum_{\ell=2}^{\infty}\sum_{j=1}^{n}\frac{(-1)^{\ell}}{p(\boldsymbol{X}_{i})^{\ell+1}}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}(\hat{p}(\boldsymbol{X}_{i})-p(\boldsymbol{X}_{i}))^{\ell}=\mathcal{O}_{L^{2}}(n^{-1}),$ (7.12)
so that
$\mathsf{E}\left[(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))^{2}\right]=\mathcal{O}(n^{-1}).$ (7.13)
@@
$\mathsf{E}\left[n^{-1}\sum_{j=2}^{\infty}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(-1)^{j}\delta_{i}}{\pi(\boldsymbol{X}_{i})^{j+1}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))^{j}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]=\mathcal{O}(n^{-1}),$ (7.14)
where the constant implicit in the error term depends on $p_{\min}$ and $\pi_{\min}$.
@@
$n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$ (7.15)
$\quad=n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)$
$\qquad+n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{1}{\pi(\boldsymbol{X}_{i})}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y).$
@@
$\mathsf{E}\left[(\delta_{i}-\pi(\boldsymbol{X}_{i}))^{2}\mid\boldsymbol{X}_{1:n}\right]=\mathsf{Var}(\delta_{i}\mid\boldsymbol{X}_{1:n})=\pi(\boldsymbol{X}_{i})(1-\pi(\boldsymbol{X}_{i})),$
@@
$\quad=\mathsf{E}\left[(\delta_{i}-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}\right]\mathsf{E}\left[(\delta_{j}-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}\right]\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}=0.$
Hence, recalling that $\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i})=n^{-1}\sum_{j=1}^{n}(\delta_{j}-\pi(\boldsymbol{X}_{i}))\,\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}/\hat{p}(\boldsymbol{X}_{i})$, we deduce
$\mathsf{E}\left[\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}\right]$
$\quad=\mathsf{E}\left[\frac{\mathsf{E}\left[(\delta_{i}-\pi(\boldsymbol{X}_{i}))^{2}\mid\boldsymbol{X}_{1:n}\right]}{n\,\pi(\boldsymbol{X}_{i})^{2}\hat{p}(\boldsymbol{X}_{i})}\right]+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\mathsf{E}\left[\frac{\mathsf{E}\big[(\delta_{i}-\pi(\boldsymbol{X}_{i}))(\delta_{j}-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{\boldsymbol{X}_{j}=\boldsymbol{X}_{i}\}}\mid\boldsymbol{X}_{1:n}\big]}{n\,\pi(\boldsymbol{X}_{i})^{2}\hat{p}(\boldsymbol{X}_{i})}\right]$
$\quad=\mathsf{E}\left[\frac{\pi(\boldsymbol{X}_{i})(1-\pi(\boldsymbol{X}_{i}))}{n\,\pi(\boldsymbol{X}_{i})^{2}\hat{p}(\boldsymbol{X}_{i})}\right]+0$
@@
$\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)\right]$
$\quad=\mathsf{E}\left[n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\mathsf{E}\left[\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})^{2}}(\hat{\pi}_{i}(\boldsymbol{X}_{1:n})-\pi(\boldsymbol{X}_{i}))\mid\boldsymbol{X}_{1:n}\right]\mathsf{E}[\mathds{1}_{\{Y_{i}\leq k/m\}}\mid\boldsymbol{X}_{1:n}]\,b_{m,k}(y)\right]$